OpenStudy (anonymous):

Find the value of g: 9g = 11. Answers Can Anyone Help Me Please ! g = 2 g = 11/9 g = -2 g = -9/11

OpenStudy (anonymous):



OpenStudy (anonymous):

g= 11/9

OpenStudy (anonymous):

once again thank you !

OpenStudy (anonymous):

9g=11 (1/9)*9g=11*(1/9) (9/9)*g=11/9 g=11/9